OpenStudy (idealist10):

How to integrate e^(-2x)/(1+e^(-x)) dx?

OpenStudy (idealist10):

\[\int\limits_{}^{}\frac{ e ^{-2x} }{ 1+e ^{-x} } dx\]

OpenStudy (idealist10):

@ganeshie8 @jabez177 @inkyvoyd @Conqueror @mathmath333

OpenStudy (irishboy123):

i'd go \(u = e^{-x}, du = - e^{-x} dx = -u ~dx \) that already looks promising so \(\int ~ \dfrac{u^2}{1 + u } . \dfrac{-1}{u} ~ du\) \(= - \int ~ \dfrac{u}{1 + u } ~ du\) \(=- \int ~ \dfrac{1 + u - 1}{1 + u } ~ du\) \(=- \int ~ 1 - \dfrac{ 1}{1 + u } ~ du\) bish bosh